# A First-look at Spatially-resolved Infrared Supernova remnants in M33 with JWST

### v1

@@ 3 Results
Here we describe the characteristics of the SNRs in the four JWST fields (Section 2), and assess their correlation with the known multiwavelength properties from published data (Section 2.2). Table 2 shows a list of the SNRs in our four fields, along with their visibility in the infrared, optical, radio and X-ray.
3.1 SNRs in the MIRI field
Of the 21 SNRs in the MIRI field, we find 3 clearly-detected SNRs – L10-039, L10-045 and L10-071 – shown in Figure 5. The snapshots in Figure 5 show the RGB images made from: (1) JWST F560W and F2100W filters together with the HST F160W filter from PHATTER, (2) PHATTER HST images from the broadband F335W, F475W and F814W filters and (3) narrowband filter images from LGGS in the lines of H, [S II] and [O III] 444We include [O III] mainly to complete the RGB images and because photoionized nebulae are richer in atoms at higher ionization states, but the presence of [O III] itself is not a strong diagnostic for distinguishing shocked from photoionized nebulae.. The SNR morphology is not always easy to distinguish from the surrounding diffuse ISM and HII regions in the emission-line and IR images. To approximately guide the eye, we also display the [S II]/H ratio image of each SNR. Regions with [S II]/H (redder pixels) are expected to be shock-heated, while [S II]/H 0.4 (bluer pixels) are more likely to be photoionized. Note that some pixels with [S II]/H are not related to SNRs, and are instead caused by poorly-subtracted stars. Diffuse ionized gas, which has low-surface brightness pixels in H, can also show [S II]/H 0.4, and will often be found at the rims of brighter HII regions (e.g. Haffner et al., 2009; Blair & Long, 1997; Long et al., 2022). We masked all pixels in the [S II] and H images that were less than 3 times the RMS noise in the image before creating the ratio image.
@@ 3.1 SNRs in the MIRI field
3.2 SNRs in the NIRCam Field
The NIRCam field (Figure 4) only has 1 clearly-detected SNR (L10-036, Figure 8) out of the 15 known SNRs in survey area (shown individually in Figure 9). SNR L10-036 is already quite unusual with its elongated, bean-shaped appearance. The shape was noted before in the low-resolution LGGS images (e.g. Long et al., 2010) as well as in HST images by Lin et al. (2020). The NIRCam images reveal that this shape is preserved even in the near-infrared. The shell-like appearance matching the full extent of the shock implies that the IR emission is produced at the forward shock interacting with the ISM. The ubiquitous emission across the optical range, as shown by three-color HST image and narrowband LGGS image in Figure 8, indicates that a signifcant fraction of the shock has become radiative, emitting bright forbidden-lines. The SNR is located about 310 pc (projected) away from the massive HII region NGC 595. We believe the bright IR emission in L10-036 and the unusual shape is most likely arising from interaction with an inhomogenous environment. The SNR is likely already ISM-dominated based on X-ray spectra (Garofali et al., 2017), and appears in projection close to a filament of 12CO(2-1)-emitting molecular gas west of the SNR (grey contours in Figure 8; the SNR is also optically brighter on this side).
The remaining SNRs in the NIRCam field do not show any noticeable extended emission that clearly traces the SNR-emitting region like L10-036 (Figure 9). Most of these objects are dominated by stars in the field. Hints of faint emission appear in cases like L10-056 and L10-060 in Figure 9 (also see Figure 15), but these emission features could also be unrelated emission from PAH and spectral lines in the surrounding ISM. Some visible emission in F444W is present in some SNRs e.g. L10-043, L10-065, but comparison with the optical images show that these are from adjacent HII regions. We conclude therefore that this remaining sample in the NIRCam field are non-detections.
@@ 3.3.3 SNR LL14-168
3.4 SNRs in the F770W-strip
The strip is shown in Figure 15, with bright emission likely dominated by the 7.7 m PAH feature based on previous Spitzer work (e.g. Smith et al., 2007; Calapa et al., 2014). The pc-resolution F770W-strip ISM shows a complex network of bubbles and filaments, similar to F770W observations in other similar star-forming galaxies (e.g. Leroy et al., 2023; Watkins et al., 2023). The emission is brightest on average towards the center of M33, and declines with distance away from it. A more detailed analysis of the PAH emission in this strip will be presented in future work. We find four SNRs located fully within the mosaic. Two SNRs L10-070 and L10-072 had partial overlap (less than half of their optical-emitting regions within the strip), so we omitted them.
Below the image of the F770W strip in Figure 15, we also show snapshots of the SNRs in the three-panel plots. The snapshots follow the same pattern as the MIRI and NIRCam snapshots to determine the emitting region in the JWST images: an F770W flux image (in MJy/sr) shown with an arcsinh stretch to balance bright and faint structures, the continuum-subtracted LGGS H, [S II] and [O III] images to show the optical-emitting nebulae in the region, and [S II]/H with the color-scale divided at 0.4 to distinguish between photoionized and shock-heated nebulae. Bluer regions denote photo-ionized gas (e.g. HII regions) and red denotes shock-heated gas or diffuse ionized gas.
A possible detection candidate in the F770W-strip is SNR L10-060, which shows a distinct patch of F770W emission within its optical-emitting region in Figure 15. This is the smallest of the five SNRs in the strip, with a diameter of 17 pc. It also appears in the NIRCam field, and is the only SNR other than SNR L10-036 with a hint of emission in the F335M filter (although the emission is challenging to make out in the glare of the un-subtracted point sources). The shape of the SNR is not entirely clear at the resolution of the LGGS image (the SNR is only 4″ across) and with all the confusing emission around, so it is not clear if the spatial extent of the optical and F770W-emitting regions are exactly the same (like e.g. L10-036 or L10-071). A high-resolution continuum-subtracted HST H or JWST Pa could confirm whether the IR and optical-emitting regions are the same, and whether the shock is producing the F770W emission.
@@ 3.5 Multi-wavelength Properties
We show the multi-wavelength properties of the SNRs in the different JWST fields in Figure 16, as derived from the optical, radio and X-ray datasets introduced in Section 2. While a multitude of information is available, we only a display a subset that is most frequently used to assess shock conditions in extragalactic SNRs — the density-sensitive [S II]:6717/6731 line ratio, the characteristic shock velocities from the full-width half-maximum (FWHM) of the H line, and the diameter estimated by Long et al. (2010) and Lee & Lee (2014) using the narrowband optical images from the LGGS survey. We use the pyneb software (Luridiana et al., 2015) to convert the [S II] doublet ratio values from optical spectroscopy (Long et al., 2018) to electron densities, assuming a plasma temperature of 104 K (we believe this temperature assumption is justified as higher temperatures would shift S to higher ionization states). For H FWHM, we also display the corresponding shock velocities in the same way as Long et al. (2018), by subtracting out in quadrature the instrumental limit of 5 Å and an assumed 20 km/s value due to thermal line broadening. We also use published 1.4 GHz VLA radio brightnesses from White et al. (2019) and Chandra-based X-ray brightnesses from Long et al. (2010), which are both sensitive to the average ambient densities and ages of the SNRs. Since these datasets are already published and available in machine-readable form in each paper, we do not separately include them in Table 2.
SNR | Infrared (JWST) | Optical | X-ray | Radio |
MIRI Field | ||||
L10-039 | Clear | ✓ | ✓ | ✓ |
L10-045 | Clear | ✓ | ✓ | ✓ |
L10-071 | Clear | ✓ | ✓ | ✓ |
L10-046 | Partial | ✓ | ✓ | ✓ |
L10-058 | Partial | ✓ | ||
L10-061 | Partial | ✓ | ✓ | ✓ |
L10-069 | Partial | ✓ | ✓ | ✓ |
L10-078 | Partial | ✓ | ✓ | ✓ |
L10-080 | Partial | ✓ | ✓ | ✓ |
L10-041 | Non-detection | ✓ | ✓ | ✓ |
L10-052 | Non-detection | ✓ | ✓ | |
L10-057 | Ambiguous | ✓ | ✓ | ✓ |
L10-064 | Ambiguous | ✓ | ✓ | ✓ |
L10-074 | Ambiguous | ✓ | ✓ | ✓ |
L10-076 | Ambiguous | ✓ | ✓ | |
L10-081 | Ambiguous | ✓ | ✓ | ✓ |
L10-083 | Ambiguous | ✓ | ✓ | |
L10-087 | Ambiguous | ✓ | ✓ | ✓ |
L10-089 | Ambiguous | ✓ | ✓ | ✓ |
LL14-075 | Ambiguous | ✓ | ✓ | ✓ |
LL14-080 | Ambiguous | ✓ | ✓ | ✓ |
NIRCam Field | ||||
L10-036 | Clear | ✓ | ✓ | ✓ |
L10-040 | Non-detection | ✓ | ✓ | ✓ |
L10-043 | Non-detection | ✓ | ✓ | ✓ |
L10-047 | Non-detection | ✓ | ✓ | ✓ |
L10-048 | Non-detection | ✓ | ||
L10-049 | Non-detection | ✓ | ✓ | ✓ |
L10-050 | Non-detection | ✓ | ✓ | |
L10-055 | Non-detection | ✓ | ✓ | ✓ |
L10-056 | Non-detection | ✓ | ✓ | ✓ |
L10-059 | Non-detection | ✓ | ✓ | ✓ |
L10-060 | Non-detection | ✓ | ✓ | |
L10-065 | Non-detection | ✓ | ✓ | ✓ |
L10-072 | Non-detection | ✓ | ||
L10-075 | Non-detection | ✓ | ✓ | |
LL14-122 | Non-detection | ✓ | ✓ |
SNR | Infrared (JWST) | Optical | Xray | Radio |
NGC 604 | ||||
L10-118 | F470N-only | ✓ | ✓ | ✓ |
L10-124 | Clear | ✓ | ✓ | ✓ |
LL14-168 | F187N-only | ✓ | ||
F770W Strip | ||||
L10-059 | Possible emission | ✓ | ✓ | ✓ |
L10-060 | Non-detection | ✓ | ✓ | |
L10-065 | Non-detection | ✓ | ✓ | ✓ |
L10-068 | Non-detection | ✓ | ✓ | ✓ |
@@
It is clear from Figure 16 that one of the distinguishing features of the cleanly detected SNRs compared to the rest of the sample is their higher than average ambient densities. The most direct evidence of this comes from the [S II]:6717/6731 ratio which is a well-known density diagnostic. The clearly detected SNRs have doublet ratios of 0.75-1.11, which correspond to electron-densities of 450-2000 cm-3. In contrast, the partial and ambiguous MIRI SNRs have values more consistent with low densities (100 cm-3) given their [S II] doublet ratios in the range of 1.3. The scatter in the [S II]-doublet values increase as one moves to lower surface brightness SNRs. Note that the [S II]-doublet-based densities refer to densities in the extended S+ recombination zone of secondary radiative shocks driven into dense cloudlets (Allen et al., 2008). While high [S II]-doublet-based densities also imply high preshock densities, estimating the latter will require more detailed comparison between the spectra and radiative shock models (e.g. as done in Dopita et al., 2018), which is outside our scope here.
@@
Aside from the low [S II] ratios, the cleanly-detected SNRs are also distinguished by their smaller sizes, large values of H FWHM, and high H, radio and X-ray surface brightnesses, compared to the rest of the sample. The smaller sizes of the cleanly-detected SNRs further corroborate the presence of high ambient densities, as do their locations close to dense IR-bright clouds in Figure 3. If we assume the SNRs are mostly radiative (given their strong [O III] emission), that means the SNRs have at least crossed their cooling radii, which for a turbulent ISM is given by , where is the average density of the turbulent ambient medium in units of cm-3 (e.g. Martizzi et al., 2016). The radii of 8-16 pc measured for these SNRs therefore implies cm-3. The high values of H FWHM indicate the presence of higher-than-average shock velocities888As discussed in Points et al. (2019), the few 100 km/s radiative shocks observed in SNRs are likely secondary shocks driven into dense clouds, as these are more likely to reach post-shock temperatures K, where the onset of significant cooling occurs. The primary blastwave is likely propagating faster by a factor , where is the cloud density, is the intercloud density, and for clouds, ., which is a common feature in optical observations of young (few 103 yr) SNRs (e.g. Fesen & Weil, 2020; Li et al., 2021). Models predict that emission from shocks at these wavelengths increase with both density and velocity (e.g. White & Long, 1991; Allen et al., 2008; Sarbadhicary et al., 2017), so the high luminosities at these wavelengths are broadly consistent with the SNRs with high ambient densities, as well as high shock velocities.
@@
Figure 16 also explains why L10-036 is the only visible SNR in the NIRCam field. Similar to the cleanly detected MIRI SNRs (L10-039, L10-045 and L10-071), the multiwavelength properties of L10-036 are quite distinct from the other SNRs as seen in Figure 16, exhibiting much faster radiative shocks (based on the H FWHM of 8.2 Å which correspond to shock speeds of about 300 km/s), propagating into high-density ambient clouds (according to the [S II]:6717/6731 of 1.02, corresponding to electron density of 658 cm-3) than the other SNRs. Consequently, L10-036 is also among the brightest in H, X-ray and radio, and smallest in diameter in our MIRI+NIRCam sample. Note that from the higher-resolution JWST and HST images, we measure the major and minor axes of the SNR as ″ (13.7 pc), and 1.92″ (7.8 pc), smaller than the previously reported diameters of 18-22 pc from lower-resolution images (Long et al., 2010).
@@
The remaining SNRs in our survey – partially-detected and ambiguous MIRI SNRs, the non-detection NIRCam SNRs, the F770W-strip SNRs, and SNRs L10-118 and LL14-168 in NGC 604 – have comparatively unremarkable properties compared to the clearly-detected ones. Their [S II]-doublet ratios cluster at the low-density limit of 1.45, velocities rarely exceeding the instrument resolution or with hints of moderate broadening (100 km/s), and diameters, radio and X-ray brightness spanning a wide range of values. Two of the SNRs have radio brightness close to the clearly-detected SNRs, but this could be due to contribution from the bright ambient background (L10-070 is near a bright HII region, while L10-060 is the near the bright central region of M33). These properties are also generally shared by the bulk M33 SNR population (not covered in our JWST survey), as seen from the distribution of gray data points in Figure 16.
3.6 Relation with progenitor masses of SNRs

### v2

@@ 3 Results
Here we describe the characteristics of the SNRs in the four JWST fields (Section 2), and assess their correlation with the known multiwavelength properties from published data (Section 2.2).
3.1 SNRs in the Southern Arm field
Of the 21 SNRs in the Southern Arm field, we find 3 clearly-detected SNRs – L10-039, L10-045 and L10-071 – shown in Figure 5. The snapshots in Figure 5 show the RGB images made from: (1) JWST F560W and F2100W filters together with the HST F160W filter from PHATTER, (2) PHATTER HST images from the broadband F335W, F475W and F814W filters and (3) narrowband filter images from LGGS in the lines of H, [S II] and [O III] 444We include [O III] mainly to complete the RGB images and because photoionized nebulae are richer in atoms at higher ionization states, but the presence of [O III] itself is not a strong diagnostic for distinguishing shocked from photoionized nebulae.. The SNR morphology is not always easy to distinguish from the surrounding diffuse ISM and HII regions in the emission-line and IR images. To approximately guide the eye, we also display the [S II]/H ratio image of each SNR. Regions with [S II]/H (redder pixels) are expected to be shock-heated, while [S II]/H 0.4 (bluer pixels) are more likely to be photoionized. Note that some pixels with [S II]/H are not related to SNRs, and are instead caused by poorly-subtracted stars. Diffuse ionized gas, which has low-surface brightness pixels in H, can also show [S II]/H 0.4, and will often be found at the rims of brighter HII regions (e.g. Haffner et al., 2009; Blair & Long, 1997; Long et al., 2022). We masked all pixels in the [S II] and H images that were less than 3 times the RMS noise in the image before creating the ratio image.
@@ 3.1 SNRs in the Southern Arm field
3.2 SNRs in the Center field
The Center field (Figure 4) only has 1 clearly-detected SNR (L10-036, Figure 8) out of the 15 known SNRs in survey area (shown individually in Figure 9). SNR L10-036 is already quite unusual with its elongated, bean-shaped appearance. The shape was noted before in the low-resolution LGGS images (e.g. Long et al., 2010) as well as in HST images by Lin et al. (2020). The NIRCam images reveal that this shape is preserved even in the near-infrared. The shell-like appearance matching the full extent of the shock implies that the IR emission is produced at the forward shock interacting with the ISM. The ubiquitous emission across the optical range, as shown by three-color HST image and narrowband LGGS image in Figure 8, indicates that a signifcant fraction of the shock has become radiative, emitting bright forbidden-lines. The SNR is located about 310 pc (projected) away from the massive HII region NGC 595. We believe the bright IR emission in L10-036 and the unusual shape is most likely arising from interaction with an inhomogenous environment. The SNR is likely already ISM-dominated based on X-ray spectra (Garofali et al., 2017), and appears in projection close to a filament of 12CO(2-1)-emitting molecular gas west of the SNR (grey contours in Figure 8; the SNR is also optically brighter on this side).
The remaining SNRs in the Center field do not show any noticeable extended emission that clearly traces the SNR-emitting region like L10-036 (Figure 9). Most of these objects are dominated by stars in the field. Hints of faint emission appear in cases like L10-056 and L10-060 in Figure 9 (also see Figure 15), but these emission features could also be unrelated emission from PAH and spectral lines in the surrounding ISM. Some visible emission in F444W is present in some SNRs e.g. L10-043, L10-065, but comparison with the optical images show that these are from adjacent HII regions. We conclude therefore that this remaining sample in the Center field are non-detections.
@@ 3.3.3 SNR LL14-168
3.4 SNRs in the North Radial Strip
The strip is shown in Figure 15, with bright emission likely dominated by the 7.7 m PAH feature based on previous Spitzer work (e.g. Smith et al., 2007; Calapa et al., 2014). The pc-resolution North Radial Strip ISM shows a complex network of bubbles and filaments, similar to F770W observations in other similar star-forming galaxies (e.g. Leroy et al., 2023; Watkins et al., 2023). The emission is brightest on average towards the center of M33, and declines with distance away from it. A more detailed analysis of the PAH emission in this strip will be presented in future work. We find four SNRs located fully within the mosaic. Two SNRs L10-070 and L10-072 had partial overlap (less than half of their optical-emitting regions within the strip), so we omitted them.
Below the image of the North Radial Strip in Figure 15, we also show snapshots of the SNRs in the three-panel plots. The snapshots follow the same pattern as the MIRI and NIRCam snapshots to determine the emitting region in the JWST images: an F770W flux image (in MJy/sr) shown with an arcsinh stretch to balance bright and faint structures, the continuum-subtracted LGGS H, [S II] and [O III] images to show the optical-emitting nebulae in the region, and [S II]/H with the color-scale divided at 0.4 to distinguish between photoionized and shock-heated nebulae. Bluer regions denote photo-ionized gas (e.g. HII regions) and red denotes shock-heated gas or diffuse ionized gas.
A possible detection candidate in the North Radial Strip is SNR L10-060, which shows a distinct patch of F770W emission within its optical-emitting region in Figure 15. This is the smallest of the five SNRs in the strip, with a diameter of 17 pc. It also appears in the Center field, and is the only SNR other than SNR L10-036 with a hint of emission in the F335M filter (although the emission is challenging to make out in the glare of the un-subtracted point sources). The shape of the SNR is not entirely clear at the resolution of the LGGS image (the SNR is only 4″ across) and with all the confusing emission around, so it is not clear if the spatial extent of the optical and F770W-emitting regions are exactly the same (like e.g. L10-036 or L10-071). A high-resolution continuum-subtracted HST H or JWST Pa could confirm whether the IR and optical-emitting regions are the same, and whether the shock is producing the F770W emission.
@@ 3.5 Multi-wavelength Properties
We show the multi-wavelength properties of the SNRs in the different JWST fields in Figure 16, as derived from the optical, radio and X-ray datasets introduced in Section 2. While a multitude of information is available, we only a display a subset that is most frequently used to assess shock conditions in extragalactic SNRs — the density-sensitive [S II]:6717/6731 line ratio, the characteristic shock velocities from the full-width half-maximum (FWHM) of the H line, and the diameter estimated by Long et al. (2010) and Lee & Lee (2014) using the narrowband optical images from the LGGS survey. We will occasionally refer to the H FWHM in this section by its equivalent velocity FWHM in km s-1 (Long et al., 2018, similar to), by subtracting out in quadrature the instrumental limit of 5 Å and an assumed 20 km/s value due to thermal line broadening. We also use published 1.4 GHz VLA radio brightnesses from White et al. (2019) and Chandra-based X-ray brightnesses from Long et al. (2010), which are both sensitive to the average ambient densities and ages of the SNRs.
@@
It is clear from Figure 16 that one of the distinguishing features of the cleanly detected SNRs compared to the rest of the sample is their higher than average ambient densities. The most direct evidence of this comes from the [S II]:6717/6731 ratio which is a well-known density diagnostic. The clearly detected SNRs have doublet ratios of 0.75-1.11, which correspond to electron-densities888We use the pyneb software (Luridiana et al., 2015) to convert the [S II] doublet ratio values from optical spectroscopy (Long et al., 2018) to electron densities for a given plasma temperature of 450-2000 cm-3, assuming a plasma temperature of 104 K (valid for the post-shock region where [S II] ionization state is abundant, Allen et al., 2008). The actual temperature of the [S II]-emitting zone may still vary depending on other shock parameters. Assuming a factor of 2 variation in temperature, e.g. for a 5103K plasma, the [S II] doublet values correspond to a lower range of densities (330-1440 cm-3), whereas a 2104K plasma temperature corresponds to a higher range of densities (540-2540 cm-3). Regardless of the temperatures, these densities are much higher than that of the partial and ambiguous MIRI SNRs, which have values more consistent with 100 cm-3 given their [S II] doublet ratios in the range of 1.3. The scatter in the [S II]-doublet values increase as one moves to lower surface brightness SNRs. Note that the [S II]-doublet-based densities refer to densities in the extended S+ recombination zone of secondary radiative shocks driven into dense cloudlets (Allen et al., 2008). While high [S II]-doublet-based densities also imply high preshock densities, estimating the latter will require more detailed comparison between the spectra and radiative shock models (e.g. as done in Dopita et al., 2018), which is outside our scope here.
@@
Aside from the low [S II] ratios, the cleanly-detected SNRs are also distinguished by their smaller sizes, large values of H FWHM, and high H, radio and X-ray surface brightnesses, compared to the rest of the sample. The smaller sizes of the cleanly-detected SNRs further corroborate the presence of high ambient densities, as do their locations close to dense IR-bright clouds in Figure 3. If we assume the SNRs are mostly radiative (given their strong [O III] emission), that means the SNRs have at least crossed their cooling radii, which for a turbulent ISM is given by , where is the average density of the turbulent ambient medium in units of cm-3 (e.g. Martizzi et al., 2016). The radii of 8-16 pc measured for these SNRs therefore implies cm-3. The high values of H FWHM indicate the presence of higher-than-average shock velocities999As discussed in Points et al. (2019), the few 100 km/s radiative shocks observed in SNRs are likely secondary shocks driven into dense clouds, as these are more likely to reach post-shock temperatures K, where the onset of significant cooling occurs. The primary blastwave is likely propagating faster by a factor , where is the cloud density, is the intercloud density, and for clouds, ., which is a common feature in optical observations of young (few 103 yr) SNRs (e.g. Fesen & Weil, 2020; Li et al., 2021). Models predict that emission from shocks at these wavelengths increase with both density and velocity (e.g. White & Long, 1991; Allen et al., 2008; Sarbadhicary et al., 2017), so the high luminosities at these wavelengths are broadly consistent with the SNRs with high ambient densities, as well as high shock velocities.
@@
Figure 16 also explains why L10-036 is the only visible SNR in the Center field. Similar to the cleanly detected MIRI SNRs (L10-039, L10-045 and L10-071), the multiwavelength properties of L10-036 are quite distinct from the other SNRs as seen in Figure 16, exhibiting much faster radiative shocks (based on the H FWHM of 8.2 Å which correspond to shock speeds of about 300 km/s), propagating into high-density ambient clouds (according to the [S II]:6717/6731 of 1.02, corresponding to electron density of 658 cm-3) than the other SNRs. Consequently, L10-036 is also among the brightest in H, X-ray and radio, and smallest in diameter in our MIRI+NIRCam sample. Note that from the higher-resolution JWST and HST images, we measure the major and minor axes of the SNR as ″ (13.7 pc), and 1.92″ (7.8 pc), smaller than the previously reported diameters of 18-22 pc from lower-resolution images (Long et al., 2010).
@@
The remaining SNRs in our survey – partially-detected and ambiguous MIRI SNRs, the non-detection NIRCam SNRs, the North Radial Strip SNRs, and SNRs L10-118 and LL14-168 in NGC 604 – have comparatively unremarkable properties compared to the clearly-detected ones. Their [S II]-doublet ratios cluster at the low-density limit of 1.45, velocities rarely exceeding the instrument resolution or with hints of moderate broadening (100 km/s), and diameters, radio and X-ray brightness spanning a wide range of values. Two of the SNRs have radio brightness close to the clearly-detected SNRs, but this could be due to contribution from the bright ambient background (L10-070 is near a bright HII region, while L10-060 is the near the bright central region of M33). These properties are also generally shared by the bulk M33 SNR population (not covered in our JWST survey), as seen from the distribution of gray data points in Figure 16.
3.6 Progenitor masses of Southern Arm SNRs